# Modeling Saliency Dataset Bias

### v1

@@ 7 Discussion
An intriguing outcome of our work is that our new model outperforms estimates of inter-observer consistency on many datasets (Sec. 5, Sec. 5). Given that our model still shows some clear failure cases (Fig. 6b), this suggests that the standard methods to estimate per-image inter-observer consistency, usually as KDE [63, 28], may be no longer be sufficient. Future efforts should consider new ways to establish a gold standard, such as combining high-performing deep neural networks (DNNs) with models of inter-observer consistency to account for consistent behavior not yet captured by DNNs and guide future model developments.
Acknowlegements
This work was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, project number: 276693517. MB further acknowledges a gift from the Open Philanthropy Foundation funded by the Good Ventures Foundation, and support from the Machine Learning Cluster of Excellence, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645.
@@ Appendix A Full results on the MIT/Tuebingen Saliency Benchmark
In Appendix A, A and A we list the full evaluation on the MIT/Tuebingen Saliency Benchmark for MIT300, CAT2000 and COCO-Freeview, including the metrics that we had to skip in the main text due to space reasons.
Table 3: MIT300 Benchmark
@@
Appendix B Bias parameter sensitivity analysis:
@@
Appendix C Generalization and adaptation on new datasets
Table 6: Kienzle Dataset
@@ Appendix C Generalization and adaptation on new datasets
The Kienzle dataset [25] consists of only 200 images which are random crops of grayscale images of natural scenes, making it an challening testcase. On this dataset, genrealization already results in substantially improved performance compared to earlier models (8% in AUC). Adapting the dataset biase parameters to the Kienzle dataset improves performance further (Appendix C). In Appendix C we also test the Toronto dataset [3].
@@
In Table C, we apply our model to the Toronto dataset [3]. The Toronto dataset consists of 120 images and hence is too small for training full deep learning models which makes it an interesting test case. On the Toronto dataset, generalization results in improved performance compared to earlier models. Adapting the few dataset parameters to the Toronto dataset improves performance further. Overall, the performance boost, however, is not as large as on the Kienzle dataset. This shows that the Toronto dataset is closer to common saliency datasets and emphasizes the need for new challenging saliency datasets.
@@
We also tested our model on the SALICON dataset. To that end, since our default models all are pretrained on SALICON, we trained a new version of the full model without previous pretraining on SALICON and then again tested generalization and adaptation, comparing to full training on SALICON (Tab. 8). We see that the achievable information gain is 0.31 bit/fix. The model trained on our combined dataset (without pretraining on SALICON) and applied with average dataset parameters performs very bad, even slightly worse than the center bias alone. However, adapting the dataset parameters to SALICON results in a performance of 0.26bit/fix, closing 85% of the generalization gap. This is in line with the results from our leave-one-dataset-out generalization test. In the case of SALICON the dataset parameters seem to account for even more of the generalization gap. This might be due to the extremely different experimental setup of SALICON (e.g., mouse traces instead of eye movements, mechanical turk instead of controlled lab environment).
Appendix D Considerations for good centerbias models
@@
For each of our five datasets and random subsets thereof, we fit the different models and evaluate on the corresponding validation splits. The results are visible in Fig. 10. We see that the first option (“KDE”) sometimes results in bad scores if little data is available. The second option (“Gaussian + uniform”) performs much better in these cases but fails to reach the performance of the nonparametric centerbias with more data. The third option (“KDE+Gaussian+uniform”) combines the advantages of both: reasonable performance already with a few images and good convergence with more data. Interestingly, whether the KDE or Gaussian+uniform performs better for low data is different from dataset to dataset. This is why we select the third option for our low-data adaptation experiments.
These results also serve to demonstrate that modeling the centerbias as a simple Gaussian is not sufficient for many datasets and can result in substantial performance penalties (see also Fig. 11).
Appendix E Multiscale ablation
We evaluated the benefits of our multiscale feature extraction state by training models with only a single scale of 1024px (which we found to be optimal across all used scales), 6 scales and 10 scales. We also varied whether the scale weights are dataset specific or dataset agnostic. The results are presented in Figure 12. We find that the multiscale backbone improves prediction performance with more scales, up to about 0.015 bit/fix.
@@ Appendix E Multiscale ablation
Appendix F The joint model excels at hard images
@@
Appendix G Details about the model architecture
@@
The scales were chosen to include 17.5 px/dva which is the scale of DeepGaze IIE (MIT1003 has 35px/dva, and DeepGaze IIE downsamples by a factor of 2). From there on we added larger scales until we ran into computational constraints, and smaller scales to the point that we still considered sensible. For the relative scales, the approach was similar: 512 pixel corresponds to the resolution that DeepGaze IIE uses internally on its training dataset, from there on we added smaller and larger scales. In an Ablation (see Appendix E), we found that including the larger scales is crucial for improving prediction performance.
@@
Appendix H Details about the datasets
@@
Appendix I Details about the training
@@
Appendix J Baseline models
@@
Appendix M Compute Ressources

### v2

@@ 7 Discussion
An intriguing outcome of our work is that our new model outperforms estimates of inter-observer consistency on many datasets (Sec.˜5, Sec.˜5). Given that our model still shows some clear failure cases (Fig.˜6b), this suggests that the standard methods to estimate per-image inter-observer consistency, usually as KDE [69, 32], may no longer be sufficient. Future efforts should consider new ways to establish a gold standard, such as combining high-performing deep neural networks (DNNs) with models of inter-observer consistency to account for consistent behavior not yet captured by DNNs and guide future model developments.
Acknowledgments
This work was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, project number: 276693517. HSK was supported by the Tübingen AI Center.
@@ Appendix A Full results on the MIT/Tuebingen Saliency Benchmark
In Sec.˜5, 5 and 5 we list the full evaluation on the MIT/Tuebingen Saliency Benchmark for MIT300, CAT2000 and COCO-Freeview, including the metrics that we had to skip in the main text due to space reasons.
@@
Appendix B Comparison with more saliency models
In Table B we extend the results table from the main paper with four additional saliency models, which are not DNN based but take inspiration from neuroscience and psychology: Itti & Koch [24] (using the implementation of [20]), RARE2012 [51], GBVS [20] and CovSal [18]. From those models, on most datasets, CovSal performs best with the exception GBVS performs substantially better. Interestingly, on CAT2000, CovSal and GBVS perform substantially better than the deep learning model EML-Net, and RARE2012 performs more similarly, but still better than EML-Net.
Model MIT1003 CAT2000 COCO Freeview DAEMONS FIGRIM average
IG AUC IG AUC IG AUC IG AUC IG AUC IG AUC
Itti & Koch - 0.757 - 0.759 - 0.702 - 0.699 - 0.766 - 0.736
RARE2012 - 0.772 - 0.777 - 0.771 - 0.706 - 0.787 - 0.762
GBVS - 0.803 - 0.802 - 0.796 - 0.710 - 0.821 - 0.786
CovSal - 0.809 - 0.847 - 0.803 - 0.679 - 0.835 - 0.795
EML-NET - 0.842 - 0.766 - 0.817 - 0.766 - 0.832 - 0.805
SalFBNet - 0.883 - 0.858 - 0.868 - 0.774 - 0.886 - 0.854
Appendix C Bias parameter sensitivity analysis:
@@
Appendix D Generalization and adaptation on new datasets
@@
The Kienzle dataset [29] consists of only 200 images which are random crops of grayscale images of natural scenes, making it an challening testcase. On this dataset, genrealization already results in substantially improved performance compared to earlier models (8% in AUC). Adapting the dataset biase parameters to the Kienzle dataset improves performance further (Appendix˜B). In Appendix D we also test the Toronto dataset [5].
@@
In Table B, we apply our model to the Toronto dataset [5]. The Toronto dataset consists of 120 images and hence is too small for training full deep learning models which makes it an interesting test case. On the Toronto dataset, generalization results in improved performance compared to earlier models. Adapting the few dataset parameters to the Toronto dataset improves performance further. Overall, the performance boost, however, is not as large as on the Kienzle dataset. This shows that the Toronto dataset is closer to common saliency datasets and emphasizes the need for new challenging saliency datasets.
@@
We also tested our model on the SALICON dataset. To that end, since our default models all are pretrained on SALICON, we trained a new version of the full model without previous pretraining on SALICON and then again tested generalization and adaptation, comparing to full training on SALICON (Tab.˜9). We see that the achievable information gain is 0.31 bit/fix. The model trained on our combined dataset (without pretraining on SALICON) and applied with average dataset parameters performs very bad, even slightly worse than the center bias alone. However, adapting the dataset parameters to SALICON results in a performance of 0.26bit/fix, closing 85% of the generalization gap. This is in line with the results from our leave-one-dataset-out generalization test. In the case of SALICON the dataset parameters seem to account for even more of the generalization gap. This might be due to the extremely different experimental setup of SALICON (e.g., mouse traces instead of eye movements, mechanical turk instead of controlled lab environment).
Refer to captionRefer to captionRefer to caption(a)(b)(c)
Appendix E Considerations for good centerbias models
@@
For each of our five datasets and random subsets thereof, we fit the different models and evaluate on the corresponding validation splits. The results are visible in Fig.˜12. We see that the first option (“KDE”) sometimes results in bad scores if little data is available. The second option (“Gaussian + uniform”) performs much better in these cases but fails to reach the performance of the nonparametric centerbias with more data. The third option (“KDE+Gaussian+uniform”) combines the advantages of both: reasonable performance already with a few images and good convergence with more data. Interestingly, whether the KDE or Gaussian+uniform performs better for low data is different from dataset to dataset. This is why we select the third option for our low-data adaptation experiments.
These results also serve to demonstrate that modeling the centerbias as a simple Gaussian is not sufficient for many datasets and can result in substantial performance penalties (see also Fig.˜13).
Appendix F Multiscale ablation
We evaluated the benefits of our multiscale feature extraction state in an ablation study where we trained the jointly trained model in different settings: we varied whether the model used absolute, relative or both scales. We also varied the numbers of scale per type and whether we added scales starting with the low or high resolutions. We find that the large scales are crucial for performance (Fig. 11a, b). We also evaluated computational demand via epoch times and find acceptable performance tradeoffs with fewer but large, preferably relative, scales, resulting in a few percent performance drop (Fig. 11c).
@@ Appendix F Multiscale ablation
Appendix G The joint model excels at hard images
@@
Appendix H Details about the model architecture
Mathematical model formulation
Given an image $I$, we denote with $R_{k,j}(I)$ the rescaled images ($k=1,2$ differentiates between the two scale type “absolute” and “relative”, and $j=1\dots 5$ indexes the specific resolutions). We extract deep features $F_{k,j}(I)=F(R_{k,j}(I))$ with our backbone $F$. Given weights $\lambda_{k,j}\geq 0$, $\sum_{k,j}\lambda_{k,j}=1$ we then compute the averaged deep features $\bar{F}=\sum_{k,j}\lambda_{k,}(R^{\prime}(F_{k,j}))$, where $R^{\prime}$ indicates a rescaling operation that rescales all deep features to the same resolution. From $\bar{F}$, the readout network $RN$ computes a spatial priority map $S=RN(\bar{F})$, which is postprocessed with the priority scaling $p$, the Gauss blur size $\sigma$, the center bias distribution $p_{\text{cb}}(x\mid I)$ and the center bias weight $\beta$ to yield the prediction $\hat{p}(x\mid I)=\text{softmax}(\mathcal{G}_{\sigma}(p\cdot S)+\beta\log p_{\text{cb}}(x\mid I))$.
@@ Appendix H Details about the model architecture
The scales were chosen to include 17.5 px/dva which is the scale of DeepGaze IIE (MIT1003 has 35px/dva, and DeepGaze IIE downsamples by a factor of 2). From there on we added larger scales until we ran into computational constraints, and smaller scales to the point that we still considered sensible. For the relative scales, the approach was similar: 512 pixel corresponds to the resolution that DeepGaze IIE uses internally on its training dataset, from there on we added smaller and larger scales. In an Ablation (see Appendix F), we found that including the larger scales is crucial for improving prediction performance.
@@
Appendix I Details about the datasets
@@
Appendix J Details about the training
@@
Appendix K Baseline models
@@
Appendix N Compute Ressources